# Whole-body Motion Planning and Safety-Critical Control for Aerial Manipulation

### v1

@@ 1 Introduction
Aerial manipulation combines the maneuverability of unmanned aerial vehicles with the dexterity of robot arms. Leveraging this synergy, numerous studies have been conducted on diverse applications, including door opening (Lee et al. (2020)), plug-pulling (Byun et al. (2021)), non-destructive testing (NDT) (D’Angelo et al. (2025)), painting (Susbielle et al. (2025)), and pick-and-place (Blöchlinger et al. (2025)).
@@ 1.1 Related Works
There have been several studies on collision avoidance of aerial manipulators. In Seo et al. (2017), the authors present a locally optimal trajectory planning method for aerial manipulation in constrained environments, using a signed distance metric and a convex mesh representation to model obstacles. In Cao et al. (2024), the authors propose a motion planning approach for aerial pick-and-place tasks that incorporates collision avoidance constraints between the movable range of the robot arm and surrounding obstacles. Additionally, Zhang et al. (2025) introduces an end-effector-oriented motion planner for aerial manipulators, which accounts for system dynamics and collision avoidance within constrained spaces. However, these methods only focus on the poses of the multirotor and the end-effector, not the collisions between the robot arm’s linkages and obstacles.
To prevent possible collisions between the robot arm’s linkages and the surrounding environment, there exist other studies that investigate collision avoidance methods that explicitly consider the interaction between these linkages and obstacles. Liu et al. (2024) presents a model predictive control framework for safe and compliant aerial manipulation, employing simple box constraints on the states of both the multirotor and the robot arm. Additionally, Lee et al. (2025) proposes a whole-body planning method for omnidirectional aerial manipulators, where both the multirotor platform and the robot arm’s linkages are represented as ellipsoids. Despite these advances, such approaches rely on bounding box or ellipsoid-based geometric representations for aerial manipulators and obstacles, which increase conservativeness when navigating through narrow passages.
To overcome the conservative issue, several studies have developed sampling-based methods for collision avoidance in aerial manipulation. In Kim et al. (2019), the authors present a rapidly-exploring random trees star (RRT*) -based planner for collision-free aerial pick-and-place tasks, and Yavari et al. (2022) introduces an RRT*-based approach that divides operational modes into approach and manipulation phases. However, sampling-based methods are generally more time-consuming than approaches based on geometric representations.
@@ 1.2 Contributions
In this paper, we propose a whole-body motion planning and a safety-critical control of an aerial manipulator using superquadrics and proxies, wherein both the multirotor platform and the robot arm’s linkages independently achieve collision avoidance with surrounding obstacles. To this end, we first derive the dynamic model of the aerial manipulator attaching a three degree-of-freedom (DOF) robot arm on a fully actuated hexarotor. Then, we design a maximum-clearance whole-body motion planner by integrating the SQ–proxy representation with Voronoi diagrams on an equilibrium manifold (Campolo and Cardin (2025)). Building upon this planner, a controller complying with propeller thrust limits and collision avoidance is constructed, where the controller framework includes a disturbance observer (DOB)-based control law presented in Back and Shim (2009). In particular, our main contributions are arranged as follows:
@@ 2 Aerial Manipulator System
According to Byun et al. (2025), the Euler-Lagrange model of the aerial manipulator is formulated as follows:
@@ 3 Whole-body motion planning
Building upon Yang et al. (2024, 2025) and Kana et al. (2021), we model the entire aerial manipulator system as well as environmental obstacles using SQs. The interactions between any two SQs are represented by proxies. Meanwhile, the end-effector of the robot arm follows Voronoi edges, while collision avoidance is also handled via proxies. The entire pipeline is illustrated in Fig. 3.
@@ 3.1 Superquadrics and proxies
SQs offer a computationally efficient representation of real-world object geometry (Jaklic et al. (2000)). We define an SQ as follows:
@@
To ensure the safety and stability of the multirotor, excessive tilting is restricted. Under this assumption, the task can be modeled in a 2D scenario. Accordingly, Eqs. 4 and 3.1 reduce to their 2D formulations, involving only the parameters $a_{1},a_{2},\epsilon,$ and $\gamma\in[-\pi,\pi]$, as presented in Yang et al. (2025).
@@ 3.2 Maximum clearance path planning
To enhance the safety of the aerial manipulator system, we adopt the Voronoi diagram as the foundation of our motion planner, since it naturally provides maximal clearance from obstacles in Arslan and Koditschek (2019). Following the approach in Yang et al. (2024), the Voronoi diagram is constructed using SQs and proxies.
@@ 3.3 Planning on equilibrium manifold
Following Yang et al. (2025); Campolo and Cardin (2025), we formulate the planning task as a manipulation potential function $W:\mathcal{Z}\times\mathcal{U}\rightarrow\mathbb{R}$, where $\boldsymbol{\mathbf{u}}$} denotes the attracting pose of the robotic end effector in control space $\mathcal{U}$, and $\boldsymbol{\mathbf{z}}$} denotes the state of the aerial manipulator system in state space $\mathcal{Z}$. The objective is to generate a smooth trajectory such that the end effector reaches the target while the entire system avoids collisions. This problem can be interpreted as finding an equilibrium state $\boldsymbol{\mathbf{z}}$$^{*}$ on the equilibrium manifold (EM) of the system:
@@
We define $\partial_{{\boldsymbol{\mathbf{q}}}}W\equiv[\partial_{q_{1}}W,\ldots,\partial_{q_{a}}W]^{T}$, where the partial operator is defined as $\partial_{{\boldsymbol{\mathbf{q}}}}=[\partial_{q_{1}},\ldots,\partial_{q_{a}}]^{T}$. Since the equilibrium state $\boldsymbol{\mathbf{z}}$$^{*}$ is implicitly defined, we adopt the adaptive ODE introduced in Yang et al. (2025) to update $\boldsymbol{\mathbf{z}}$$^{*}$, where $\eta$ represents step size in the ODE :
@@ 4.1 Inner-Loop Control Law
To cope with three performance deteriorating factors, i.e., movement of the attached robot arm, model uncertainties, and time-varying external disturbance, we adopt a DOB-based controller extending the control law introduced in Kim et al. (2017) to incorporate 6 DOF pose.
@@
$\begin{split}\boldsymbol{T}=&\boldsymbol{B^{-1}}(\boldsymbol{\phi})\{\boldsymbol{\hat{M}}(\boldsymbol{\phi})(\boldsymbol{K_{d}}\boldsymbol{\dot{e}}+\boldsymbol{K_{p}}\boldsymbol{e})\\ &\qquad\qquad\qquad+\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{\hat{G}}-\boldsymbol{\hat{d}}\}\\ \boldsymbol{\hat{d}}=&\boldsymbol{E_{1}}\boldsymbol{p_{dob}}-\boldsymbol{E_{2}}\boldsymbol{\dot{q}_{dob}}+\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{\hat{G}}\\ \boldsymbol{\dot{q}_{dob}}=&\boldsymbol{A_{dob}}\boldsymbol{q_{dob}}+\boldsymbol{B_{dob}}\boldsymbol{q}\\ \boldsymbol{\dot{p}_{dob}}=&\boldsymbol{A_{dob}}\boldsymbol{p_{dob}}+\boldsymbol{B_{dob}}\boldsymbol{M^{-1}(\boldsymbol{\phi})}\boldsymbol{B}(\boldsymbol{\phi})\boldsymbol{T}\end{split}$ (15)
@@ 4.3 Outer-Loop Control Law
Meanwhile, since $h_{co}$ is the function of $(\boldsymbol{q},\boldsymbol{\theta})$, we need to differentiate it twice to attain the thrust inputs. Hence, Eq. 18 becomes a high-order CBF (Xiao and Belta (2021)), and the collision avoidance condition is formulated as follows:
@@ 5.2 Benchmark
We compare our method against two widely used sampling-based algorithms: RRT* (Karaman and Frazzoli (2011)) and Model Predictive Path Integral control (MPPI) (Williams et al. (2017)), as well as an ablation study that replaces SQs with ellipses for obstacle modeling. The comparisons are illustrated in Fig. 6 and Table 1. In the implementation, RRT* is sampled in ${\boldsymbol{\mathbf{z}}}_{sys}$, with collision detection based on the geometric overlap between the aerial manipulator and obstacles at each configuration $\boldsymbol{\mathbf{z}}$}. MPPI is evaluated with a cost function that penalizes both target distance and trajectory smoothness, using N = 50 rollouts per control step, which provides a trade-off between computational cost and trajectory quality. The ablation study adopts the same settings as the proposed algorithm, except that obstacles are represented as ellipses instead of SQs.

### v2

@@ 1 Introduction
Aerial manipulation combines the maneuverability of unmanned aerial vehicles with the dexterity of robot arms. Leveraging this synergy, numerous studies have been conducted on diverse applications, including door opening (lee2020aerial), plug-pulling (byun2021stability), non-destructive testing (NDT) (d2025semi), painting (susbielle2025pointillism), and pick-and-place (blochlinger2025reliable).
@@ 1.1 Related Works
There have been several studies on collision avoidance of aerial manipulators. In seo2017locally, the authors present a locally optimal trajectory planning method for aerial manipulation in constrained environments, using a signed distance metric and a convex mesh representation to model obstacles. In cao2024motion, the authors propose a motion planning approach for aerial pick-and-place tasks that incorporates collision avoidance constraints between the movable range of the robot arm and surrounding obstacles. Additionally, zhang2025end introduces an end-effector-oriented motion planner for aerial manipulators, which accounts for system dynamics and collision avoidance within constrained spaces. However, these methods only focus on the poses of the multirotor and the end-effector, not the collisions between the robot arm’s linkages and obstacles.
To prevent possible collisions between the robot arm’s linkages and the surrounding environment, there exist other studies that investigate collision avoidance methods that explicitly consider the interaction between these linkages and obstacles. liu2024coordinated presents a model predictive control framework for safe and compliant aerial manipulation, employing simple box constraints on the states of both the multirotor and the robot arm. Additionally, lee2025autonomous proposes a whole-body planning method for omnidirectional aerial manipulators, where both the multirotor platform and the robot arm’s linkages are represented as ellipsoids. Despite these advances, such approaches rely on bounding box or ellipsoid-based geometric representations for aerial manipulators and obstacles, which increase conservativeness when navigating through narrow passages.
To overcome the conservative issue, several studies have developed sampling-based methods for collision avoidance in aerial manipulation. In kim2019sampling, the authors present a rapidly-exploring random trees star (RRT*) -based planner for collision-free aerial pick-and-place tasks, and yavari2022interleaved introduces an RRT*-based approach that divides operational modes into approach and manipulation phases. However, sampling-based methods are generally more time-consuming than approaches based on geometric representations.
@@ 1.2 Contributions
In this paper, we propose a whole-body motion planning and a safety-critical control of an aerial manipulator using superquadrics and proxies, wherein both the multirotor platform and the robot arm’s linkages independently achieve collision avoidance with surrounding obstacles. To this end, we first derive the dynamic model of the aerial manipulator attaching a three degree-of-freedom (DOF) robot arm on a fully actuated hexarotor. Then, we design a maximum-clearance whole-body motion planner by integrating the SQ–proxy representation with Voronoi diagrams on an equilibrium manifold (campolo2025geometric). Building upon this planner, a controller complying with propeller thrust limits and collision avoidance is constructed, where the controller framework includes a disturbance observer (DOB)-based control law presented in back2009inner. In particular, our main contributions are arranged as follows:
@@ 2 Aerial Manipulator System
According to byun2025safety, the Euler-Lagrange model of the aerial manipulator is formulated as follows:
@@ 3 Whole-body motion planning
Building upon yang2024path; yang2025planning and kana2021human, we model the entire aerial manipulator system as well as environmental obstacles using SQs. The interactions between any two SQs are represented by proxies. Meanwhile, the end-effector of the robot arm follows Voronoi edges, while collision avoidance is also handled via proxies. The entire pipeline is illustrated in Fig. 3.
@@ 3.1 Superquadrics and proxies
SQs offer a computationally efficient representation of real-world object geometry (jaklic2000segmentation). We define an SQ as follows:
@@
To ensure the safety and stability of the multirotor, excessive tilting is restricted. Under this assumption, the task can be modeled in a 2D scenario. Accordingly, Eqs. 4 and 3.1 reduce to their 2D formulations, involving only the parameters $a_{1},a_{2},\epsilon,$ and $\gamma\in[-\pi,\pi]$, as presented in yang2025planning.
@@ 3.2 Maximum clearance path planning
To enhance the safety of the aerial manipulator system, we adopt the Voronoi diagram as the foundation of our motion planner, since it naturally provides maximal clearance from obstacles in arslan2019sensor. Following the approach in yang2024path, the Voronoi diagram is constructed using SQs and proxies.
@@ 3.3 Planning on equilibrium manifold
Following yang2025planning; campolo2025geometric, we formulate the planning task as a manipulation potential function $W:\mathcal{Z}\times\mathcal{U}\rightarrow\mathbb{R}$, where $\boldsymbol{\mathbf{u}}$} denotes the attracting pose of the robotic end effector in control space $\mathcal{U}$, and $\boldsymbol{\mathbf{z}}$} denotes the state of the aerial manipulator system in state space $\mathcal{Z}$. The objective is to generate a smooth trajectory such that the end effector reaches the target while the entire system avoids collisions. This problem can be interpreted as finding an equilibrium state $\boldsymbol{\mathbf{z}}$$^{*}$ on the equilibrium manifold (EM) of the system:
@@
We define $\partial_{{\boldsymbol{\mathbf{q}}}}W\equiv[\partial_{q_{1}}W,\ldots,\partial_{q_{a}}W]^{T}$, where the partial operator is defined as $\partial_{{\boldsymbol{\mathbf{q}}}}=[\partial_{q_{1}},\ldots,\partial_{q_{a}}]^{T}$. Since the equilibrium state $\boldsymbol{\mathbf{z}}$$^{*}$ is implicitly defined, we adopt the adaptive ODE introduced in yang2025planning to update $\boldsymbol{\mathbf{z}}$$^{*}$, where $\eta$ represents step size in the ODE :
@@ 4.1 Inner-Loop Control Law
To cope with three performance deteriorating factors, i.e., movement of the attached robot arm, model uncertainties, and time-varying external disturbance, we adopt a DOB-based controller extending the control law introduced in kim2017robust to incorporate 6 DOF pose.
@@
$\begin{split}\boldsymbol{T}=&\boldsymbol{B^{-1}}(\boldsymbol{\phi})\{\boldsymbol{\hat{M}}(\boldsymbol{\phi})(\boldsymbol{K_{d}}\boldsymbol{\dot{e}}+\boldsymbol{K_{p}}\boldsymbol{e})\\ &\qquad\qquad\qquad+\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{\hat{G}}-\boldsymbol{\hat{d}}\}\\ \boldsymbol{\hat{d}}=&-\boldsymbol{\hat{M}}(\boldsymbol{\phi})(\boldsymbol{E_{1}}\boldsymbol{p_{dob}}-\boldsymbol{E_{2}}\boldsymbol{\dot{q}_{dob}})+\boldsymbol{\hat{C}}(\boldsymbol{\phi},\boldsymbol{\dot{\phi}})+\boldsymbol{\hat{G}}\\ \boldsymbol{\dot{q}_{dob}}=&\boldsymbol{A_{dob}}\boldsymbol{q_{dob}}+\boldsymbol{B_{dob}}\boldsymbol{q}\\ \boldsymbol{\dot{p}_{dob}}=&\boldsymbol{A_{dob}}\boldsymbol{p_{dob}}+\boldsymbol{B_{dob}}\boldsymbol{M^{-1}(\boldsymbol{\phi})}\boldsymbol{B}(\boldsymbol{\phi})\boldsymbol{T}\end{split}$ (15)
@@ 4.3 Outer-Loop Control Law
Meanwhile, since $h_{co}$ is the function of $(\boldsymbol{q},\boldsymbol{\theta})$, we need to differentiate it twice to attain the thrust inputs. Hence, Eq. 18 becomes a high-order CBF (xiao2021high), and the collision avoidance condition is formulated as follows:
@@ 5.2 Benchmark
We compare our method against two widely used sampling-based algorithms: RRT* (karaman2011sampling) and Model Predictive Path Integral control (MPPI) (williams2017information), as well as an ablation study that replaces SQs with ellipses for obstacle modeling. The comparisons are illustrated in Fig. 6 and Table 1. In the implementation, RRT* is sampled in ${\boldsymbol{\mathbf{z}}}_{sys}$, with collision detection based on the geometric overlap between the aerial manipulator and obstacles at each configuration $\boldsymbol{\mathbf{z}}$}. MPPI is evaluated with a cost function that penalizes both target distance and trajectory smoothness, using N = 50 rollouts per control step, which provides a trade-off between computational cost and trajectory quality. The ablation study adopts the same settings as the proposed algorithm, except that obstacles are represented as ellipses instead of SQs.